10. Determine all points where the function /(z) = 23 + 2z+1 - is continuous 23_1
Answer #1

Criven that f(3) = 23723 +1. f(3) = 3+1 em we know that every polynomial is continuous in everywhere. (00,-). 23+23+1 continu